OpenStudy (anonymous):

helpppp!!!!!!!!!!!!!!!!!!!!! Mass box A = 10 grams; Mass box B = 5 grams; Mass box C—made of one A and one B How many boxes of A would be required to make 30 grams of C? ________

OpenStudy (anonymous):

2 boxes of A

OpenStudy (anonymous):

how did u get that answer????

OpenStudy (anonymous):

C=A+B we know C has to be 30 so 30=A+B i started by adding A and B which is 15 15 is only half of what we need so i multiplied it by 2 30=20+10 30=30 so you would need 2 boxes of each

OpenStudy (anonymous):

ok thank u....

OpenStudy (anonymous):

yeah my pleasure

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Mass box A = 10 grams; Mass box B = 5 grams; Mass box C—made of one A and one B How many grams of A would be needed to give the same number of particles: as 7.5 grams of B? 1015205 as 2.5 grams? 1057.512.5 as 12.5 grams? 5037.52512.5 as 1 gram